OpenStudy (anonymous):

need help: Consider the function f(x) = x^2-16/x+3. Determine all asymptotes of this function including horizontal, vertical, and oblique (slant). please check: What are the vertical asymptotes of the function f(x) = 5x+1/x^2+x-6? --> x can not = 2, x = -3 correct? What are the zeros of the function f(x) = 3x^2-3x-6/3x-6? --> -2 correct? Consider the function f(x) = x^2+x-12/x-3. Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph. --> disc = -4? x int = (3,-4)? y int = 3? basic shape = straight line? correct?

OpenStudy (anonymous):

is it some assignment thingy ?

OpenStudy (anonymous):

it was, i got them wrong but i know they will be on the test

OpenStudy (anonymous):

ok, for first part f(x) = x^2-16/x+3. in order to get vertical asymptotic set denominator equal to zero and solve for x. x+3=0 so can you tell me what is vertical asymptote ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

great ! thats correct

OpenStudy (anonymous):

now in order to find the Horizontal asymptote , check the following 1)if the degree of the highest variable in numerator and denominator is same then asymptote is simply the ratio of the coefficients of the largest value For example f(x)= (3x^2+8)/(5x^3+9) since both numerator and denominator have same degree so Horizontal asymtote is the ratios of their coefficients y=3/5 is horizontal asymptote.. 2) if the degree of numerator is greater than denominator, than there is no Horizontal asymptote For example f(x)= (5x^3+7)/(3x^2+8) since the degree of numerator is greater ( it is 3 x^3) and that of denominator is less ( it is 2 x^2 ) there is no horizontal asymptote. can you now tell me Horizontal asymptote of f(x) = x^2-16/x+3. ???

OpenStudy (anonymous):

hmmm (x+4)(x-4)/x+3 id say the denom is lower than num

OpenStudy (anonymous):

so none

OpenStudy (anonymous):

yes you are right. there was no need of factorization !!!. no it is not one since the degree of denominator is less so there is no Horizontal asymptotic at all.

OpenStudy (anonymous):

i c

OpenStudy (anonymous):

when there is no Horizontal asymptote then there is oblique (slant) asymtote .

OpenStudy (anonymous):

good deal

OpenStudy (anonymous):

sorry divide by x+3 for that you have to perform the Long division. so divide the (x2-16) by (x+3) and tell me what you get after division.

OpenStudy (anonymous):

x+3 / x^2 - 16 x^3-16x +3x^2-48 - x^2-16 x^3-16x +2x-32

OpenStudy (anonymous):

what is quotient ?

OpenStudy (anonymous):

x^2 - 16

OpenStudy (anonymous):

i think you went wrong !!!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

click on show steps

OpenStudy (anonymous):

very good ty

OpenStudy (anonymous):

so -7

OpenStudy (anonymous):

-7 is remainder !!! quotient is (x-3)

OpenStudy (anonymous):

i understand now

OpenStudy (anonymous):

so what ever you get in quotient set that equal to y hence y=x-3 should be slant asymptotic.

OpenStudy (anonymous):

y=3 ?

OpenStudy (anonymous):

no y=x-3 is the slant asymptotic

OpenStudy (anonymous):

so va = -3, ha = 7/(x-3), oa = y= x-3

OpenStudy (anonymous):

there is no Horizontal asymptote !!! told you already both slant and Horizontal cannot occur at the same time.

OpenStudy (anonymous):

yes that was correct

OpenStudy (anonymous):

so asymptotic stuff is now clear ??

OpenStudy (anonymous):

believe so

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

this link might be helpful . http://www.purplemath.com/modules/asymtote4.htm

OpenStudy (anonymous):

What are the vertical asymptotes of the function f(x) = 5x+1/x^2+x-6? --> x can not = 2, x = -3 correct? What are the zeros of the function f(x) = 3x^2-3x-6/3x-6? --> -2 correct? Consider the function f(x) = x^2+x-12/x-3. Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph. --> disc = -4? x int = (3,-4)? y int = 3? basic shape = straight line? correct?

OpenStudy (anonymous):

are those correct

OpenStudy (anonymous):

yes the first one is correct f(x) = 5x+1/x^2+x-6 , x=-3, x=2 they are va.

OpenStudy (anonymous):

phew

OpenStudy (anonymous):

What are the zeros of the function f(x) = 3x^2-3x-6/3x-6? --> -2 correct?

OpenStudy (anonymous):

in order to get zeros of polynomial set f(x)=0 so you need to solve \[\Large \frac{3x^2-3x-6}{3x-6}=0\] \[\Large \implies (3x^2-3x-6)=0\] solve this Quadratic equation now.

OpenStudy (anonymous):

can you do this now ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x(x^2-x-2) = x(x+1)(x-2)

OpenStudy (anonymous):

ops 3(x+1)(x-2)

OpenStudy (anonymous):

yes solve 3(x+1)=0 (x-2)=0

OpenStudy (anonymous):

zero f (x) = -1

OpenStudy (anonymous):

i put 2 originally and got it wrong

OpenStudy (anonymous):

there are two zeros because we have solved Quadratic equation. they are 3(x+1)=0 x-2=0 so x=-1 x=2 are two zeros.

OpenStudy (anonymous):

ty